OpenStudy (jiteshmeghwal9):

A merchant mixes two varieties of rice , one costing Rs.35per kg. & another at Rs.45per kg. in the ratio of 3:2 . If he sells the mixed variety at rs 41.60 per kg.,find his gain or loss percent.

OpenStudy (jiteshmeghwal9):

@mayankdevnani plz help :)

OpenStudy (jiteshmeghwal9):

@hartnn plz help :)

OpenStudy (jiteshmeghwal9):

@hartnn

OpenStudy (sirm3d):

\[\frac{ 3(35)+2(45) }{ 5 }=39\] that should be the fair price for the mixed rice

hartnn (hartnn):

yes, that ^

OpenStudy (mayankdevnani):

total cost =

OpenStudy (mayankdevnani):

35+45=Rs.80

OpenStudy (mayankdevnani):